1.the salaries of ayanda,refilwe and danisile are in the ratio 2:3:5.if ayanda's salary is R15000, find danisile's salary.
2.store A is selling 10 rolls of toilet paper for R35.50,while store B is selling 32 rolls for R98.00.write down each unit rate to the nearest cent and indicate the one that is the better deal.
1
Expert's answer
2020-10-13T19:20:41-0400

1.

Let "x=" salary of Refilwe, then we get a proportion:

"2:3=15000:x"

"\frac{2}{3}=\frac{15000}{x}"

"2\u22c5x=15000\u22c53"

"2x=45000"

"x=45000:2"

"x=22500"

Let "y=" salary of Danisile, then

"3:5=22500:y"

"\frac{3}{5}=\frac{22500}{y}"

"3\u22c5y=22500\u22c55"

"3y=112500"

"y=112500:3"

"y=37500"

Solution: The salary of Danisile is R37500.

2.

Store A:

"Rate=\frac{R35.50}{10rolls}=\frac{R3.55}{1roll}=R3.55" per roll

Store B:

"Rate=\frac{R98.00}{32roll}=\frac{3.06}{1roll}=R3.06" per roll

Unit rate of store B is the best deal.

Solution: R3.55 per roll and R3.06 per roll (the best deal).